# MaskControl: Spatio-Temporal Control for Masked Motion Synthesis

### v3

@@ 3.4 Differentiable Expectation Sampling (DES)
[…]
@@ 5 Experiment
Method Joint R-Precision Top-3 $\uparrow$ FID $\downarrow$ Diversity $\rightarrow$ Foot Skating Ratio $\downarrow$ Trajectory Error (>50 cm) () $\downarrow$ Location Error (>50 cm) () $\downarrow$ Average Error (cm) $\downarrow$
@@ 5.3 Body Part Editing
Method R-precision $\uparrow$ FID $\downarrow$ MM-Dist $\downarrow$ Diversity $\rightarrow$
@@ A.3 Pseudo Code of MaskControl Inference
1:Masked Motion Model ($MMM$), Logits Regularizer ($LR$), mask scheduling function $\gamma(\cdot)$, spatial control signals $s$ (if any), text prompts $W$ (if any).
@@
3:for all $t$ from 111 to $T$ do $\triangleright$ Unmask process in $T$ steps
4:  $\{\mathbf{f}\}\leftarrow LR(X_{\overline{\mathbf{M}}},W,s;\phi)$triangleright Logits Regularizer
5:  $l\leftarrow MMM(X_{\overline{\mathbf{M}}},\mathbf{p},\{\mathbf{f}\};\theta)$triangleright Masked Motion Model
6:  for all $i$ from 111 to $I_{l}$ do $\triangleright$ Logits Optimization
@@
9:  $X_{\overline{\mathbf{M}}}\leftarrow\gamma(l,t)$triangleright mask out tokens based on logits $l$ at time step $t$
@@
11:for all $i$ from 111 to $I_{e}$ do $\triangleright$ Logits Optimization
@@ A.4 Implementation Details
We modified the MoMask [16] model by retraining it with a cross-entropy loss applied to all tokens, instead of just the masked positions. This retrained model serves as our pretrained base model, and we kept the default hyperparameter settings unchanged. To improve robustness to text variation, we randomly drop 10% of the text conditioning, which also allows the model to be used for Classifier-Free Guidance (CFG). The weight for Eq. 5 is set to $\alpha=0.1$. We use a codebook of size 512, with embeddings of size 512 and 6 residual layers. The Transformer embedding size is set to 384, with 6 attention heads, each with an embedding size of 64, distributed across 8 layers. This configuration demonstrates the feasibility of converting between two different embedding sizes and spaces using the Differentiable Expectation Sampling (DES). The encoder and decoder downsample the motion sequence length by a factor of 4 when mapping to token space. The learning rate follows a linear warm-up schedule, reaching 2e-4 after 2000 iterations, using AdamW optimization. The mini-batch size is set to 512 for training RVQ-VAE and 64 for training the Transformers. During inference, the CFG scale is set to $cfg=4$ for the base layer and $cfg=5$ for the 6 layers of residual, with 10 steps for generation. We use pretrained CLIP model [41] to generate text embeddings, which have a size of 512. These embeddings are then projected down to a size of 384 to match the token size used by the Transformer. Logits Regularizer is a trainable copy of Masked Transformer with the zero linear layer connect to the output each layer of the Masked Transformer. During inference, Logits Optimization applies L2 loss with a learning rate of 0.060.060.06 for 100 iterations in Logits Optimization for each of the 10 generation steps and 600 iterations in the last unmasking step. We apply temperature of 1 for all 10 steps and 1e-8 for residual layers. We follow the implementation from [21, 55, 51], applying the spatial control signal only to joint positions and omitting rotations.

### v4

@@ 3.4 Differentiable Expectation Sampling (DES)
where $l$ is logits, $\tau$ refers to temperature, and $g$ represents Gumbel noise with $g_{1},\dots,g_{K}$ being independent and identically distributed (i.i.d.) samples from a $\text{Gumbel}(0,1)$ distribution. The $\text{Gumbel}(0,1)$ distribution can be sampled via inverse transform sampling by first drawing $u\sim\text{Uniform}(0,1)$ and then computing $g=-\log(-\log(u))$.
@@ 5 Experiment
Method Joint R-Precision Top-3 $\uparrow$ FID $\downarrow$ Diversity $\rightarrow$ Foot Skating Ratio $\downarrow$ Trajectory Error (>50 cm) (% ) $\downarrow$ Location Error (>50 cm) (% ) $\downarrow$ Average Error (cm) $\downarrow$
@@ 5.3 Body Part Editing
Method R-precision $\uparrow$ FID $\downarrow$ $M$-Dist $\downarrow$ Diversity $\rightarrow$
@@ A.3 Pseudo Code of MaskControl Inference
1:Masked Motion Model (MMMMMM), Logits Regularizer (LRLR), mask scheduling function $\gamma(\cdot)$, spatial control signals $s$ (if any), text prompts $W$ (if any).
@@
3:for all $t$ from 11 to $T$ do $\triangleright$ Unmask process in $T$ steps
4:  $\{\mathbf{f}\}\leftarrow LR(X_{\overline{\mathbf{M}}},W,s;\phi)$ $\triangleright$ Logits Regularizer
5:  $l\leftarrow MMM(X_{\overline{\mathbf{M}}},W,\{\mathbf{f}\};\theta)$ $\triangleright$ Masked Motion Model
6:  for all $i$ from 11 to $I_{l}$ do $\triangleright$ Logits Optimization
@@
9:  $X_{\overline{\mathbf{M}}}\leftarrow\gamma(l,t)$ $\triangleright$ mask out tokens based on logits $l$ at time step $t$
@@
11:for all $i$ from 11 to $I_{e}$ do $\triangleright$ Logits Optimization
@@ A.4 Implementation Details
We modified the MoMask [16] model by retraining it with a cross-entropy loss applied to all tokens, instead of just the masked positions. This retrained model serves as our pretrained base model, and we kept the default hyperparameter settings unchanged. To improve robustness to text variation, we randomly drop 10% of the text conditioning, which also allows the model to be used for Classifier-Free Guidance (CFG). The weight for Eq. 5 is set to $\alpha=0.1$. We use a codebook of size 512, with embeddings of size 512 and 6 residual layers. The Transformer embedding size is set to 384, with 6 attention heads, each with an embedding size of 64, distributed across 8 layers. This configuration demonstrates the feasibility of converting between two different embedding sizes and spaces using the Differentiable Expectation Sampling (DES). The encoder and decoder downsample the motion sequence length by a factor of 4 when mapping to token space. The learning rate follows a linear warm-up schedule, reaching 2e-4 after 2000 iterations, using AdamW optimization. The mini-batch size is set to 512 for training RVQ-VAE and 64 for training the Transformers. During inference, the CFG scale is set to $cfg=4$ for the base layer and $cfg=5$ for the 6 layers of residual, with 10 steps for generation. We use pretrained CLIP model [41] to generate text embeddings, which have a size of 512. These embeddings are then projected down to a size of 384 to match the token size used by the Transformer. Logits Regularizer is a trainable copy of Masked Transformer with the zero linear layer connect to the output each layer of the Masked Transformer. During inference, Logits Optimization applies L2 loss with a learning rate of $0.06$ for 100 iterations in Logits Optimization for each of the 10 generation steps and 600 iterations in the last unmasking step. We apply temperature of 1 for all 10 steps and 1e-8 for residual layers. We follow the implementation from [21, 55, 51], applying the spatial control signal only to joint positions and omitting rotations.